# Primordial Black Holes Formed during Magneto-Hydrodynamic Turbulence in Early Universe as Dominant Part of Dark Matter

### v2

@@ Acknowledgements.
- Carr [1975] B. J. Carr, The primordial black hole mass spectrum (1975).
@@
- Musco [2019] I. Musco, Threshold for primordial black holes: Dependence on the shape of the cosmological perturbations, Physical Review D 100, 123524 (2019).
@@
- Barrau [2000] A. Barrau, Primordial black holes as a source of extremely high energy cosmic rays, Astroparticle Physics 12, 269 (2000).
- Garriga et al. [2016] J. Garriga, A. Vilenkin, and J. Zhang, Black holes and the multiverse, Journal of Cosmology and Astroparticle Physics 2016 (02), 064.
- Deng et al. [2017] H. Deng, J. Garriga, and A. Vilenkin, Primordial black hole and wormhole formation by domain walls, Journal of Cosmology and Astroparticle Physics 2017 (04), 050.
- Garriga and Vilenkin [1993] J. Garriga and A. Vilenkin, Black holes from nucleating strings, Physical Review D 47, 3265 (1993).
- Yang et al. [2015] H. Yang, A. Zimmerman, and L. Lehner, Turbulent black holes, Physical review letters 114, 081101 (2015).
@@
- Okele and Rees [1980] P. Okele and M. Rees, Observational consequences of positron production by evaporating black holes, Astronomy and Astrophysics, vol. 81, no. 1-2, Jan. 1980, p. 263, 264. 81, 263 (1980).
- Belotsky and Kirillov [2015] K. Belotsky and A. Kirillov, Primordial black holes with mass 1016- 1017 g and reionization of the universe, Journal of Cosmology and Astroparticle Physics 2015 (01), 041.
- Cline et al. [1997] D. B. Cline, D. A. Sanders, and W. Hong, Further evidence for some gamma-ray bursts consistent with primordial black hole evaporation, The Astrophysical Journal 486, 169 (1997).
- Hawkins [1993] M. Hawkins, Gravitational microlensing, quasar variability and missing matter, Nature 366, 242 (1993).
- Lacey and Ostriker [1985] C. G. Lacey and J. P. Ostriker, Massive black holes in galactic halos?, Astrophysical Journal, Part 1 (ISSN 0004-637X), vol. 299, Dec. 15, 1985, p. 633-652. 299, 633 (1985).
- Yokoyama [1995] J. Yokoyama, Formation of macho-primordial black holes in inflationary cosmology, arXiv preprint astro-ph/9509027 (1995).
- Bean and Magueijo [2002] R. Bean and J. Magueijo, Could supermassive black holes be quintessential primordial black holes?, Physical Review D 66, 063505 (2002).
- Afshordi et al. [2003] N. Afshordi, P. McDonald, and D. Spergel, Primordial black holes as dark matter: The power spectrum and evaporation of early structures, The Astrophysical Journal 594, L71 (2003).
- Green et al. [2004] A. M. Green, A. R. Liddle, K. A. Malik, and M. Sasaki, New calculation of the mass fraction of primordial black holes, Physical Review D 70, 041502 (2004).

### v3

@@ Acknowledgements.
- Cai et al. [2018] Y.-F. Cai, X. Tong, D.-G. Wang, and S.-F. Yan, Primordial black holes from sound speed resonance during inflation, Physical Review Letters 121, 081306 (2018).
- Khlopov et al. [1985] M. Y. Khlopov, B. Malomed, and Y. B. Zeldovich, Gravitational instability of scalar fields and formation of primordial black holes, Monthly Notices of the Royal Astronomical Society 215, 575 (1985).
- Saini and Stojkovic [2018] A. Saini and D. Stojkovic, Modified hoop conjecture in expanding spacetimes and primordial black hole production in frw universe, Journal of Cosmology and Astroparticle Physics 2018 (05), 071.
- Barrow and Carr [1996] J. D. Barrow and B. J. Carr, Formation and evaporation of primordial black holes in scalar-tensor gravity theories, Physical Review D 54, 3920 (1996).
- Kawai and Kim [2021] S. Kawai and J. Kim, Primordial black holes from gauss-bonnet-corrected single field inflation, Physical Review D 104, 083545 (2021).
- Hawking et al. [1982] S. W. Hawking, I. G. Moss, and J. M. Stewart, Bubble collisions in the very early universe, Physical Review D 26, 2681 (1982).
- Liu et al. [2022] J. Liu, L. Bian, R.-G. Cai, Z.-K. Guo, and S.-J. Wang, Primordial black hole production during first-order phase transitions, Physical Review D 105, L021303 (2022).
- Jung and Okui [2024] T. H. Jung and T. Okui, Primordial black holes from bubble collisions during a first-order phase transition, Physical Review D 110, 115014 (2024).
@@
- Roper Pol et al. [2020] A. Roper Pol, S. Mandal, A. Brandenburg, T. Kahniashvili, and A. Kosowsky, Numerical simulations of gravitational waves from early-universe turbulence, Physical Review D 102, 083512 (2020).
- Brandenburg et al. [1996] A. Brandenburg, K. Enqvist, and P. Olesen, Large-scale magnetic fields from hydromagnetic turbulence in the very early universe, Physical Review D 54, 1291 (1996).
- Pen and Turok [2016] U.-L. Pen and N. Turok, Shocks in the early universe, Physical Review Letters 117, 131301 (2016).
@@
- Khlopov et al. [1999] M. Y. Khlopov, R. Konoplich, S. Rubin, and A. S. Sakharov, First order phase transitions as a source of black holes in the early universe, arXiv preprint hep-ph/9912422 (1999).
- Lewicki et al. [2023] M. Lewicki, P. Toczek, and V. Vaskonen, Primordial black holes from strong first-order phase transitions, Journal of High Energy Physics 2023, 1 (2023).
- Baker et al. [2021] M. J. Baker, M. Breitbach, J. Kopp, and L. Mittnacht, Primordial black holes from first-order cosmological phase transitions, arXiv preprint arXiv:2105.07481 (2021).
- Carr [1975] B. J. Carr, The primordial black hole mass spectrum (1975).
@@
- Cai and Wang [2018] R.-G. Cai and S.-J. Wang, Energy budget of cosmological first-order phase transition in flrwbackground, Science China Physics, Mechanics & Astronomy 61, 080411 (2018).
@@
- Carr et al. [2021] B. Carr, K. Kohri, Y. Sendouda, and J. Yokoyama, Constraints on primordial black holes, Reports on Progress in Physics 84, 116902 (2021).
- Niikura et al. [2019] H. Niikura, M. Takada, N. Yasuda, R. H. Lupton, T. Sumi, S. More, T. Kurita, S. Sugiyama, A. More, M. Oguri, et al., Microlensing constraints on primordial black holes with subaru/hsc andromeda observations, Nature Astronomy 3, 524 (2019).
- Micha and Tkachev [2003] R. Micha and I. I. Tkachev, Relativistic turbulence: A long way from preheating to equilibrium, Physical review letters 90, 121301 (2003).
@@
- Khachatryan [2008] V. Khachatryan, Modified kolmogorov wave turbulence in qcd matched onto “bottom-up” thermalization, Nuclear Physics A 810, 109 (2008).
- Safarzadeh [2018] M. Safarzadeh, Primordial black holes as seeds of magnetic fields in the universe, Monthly Notices of the Royal Astronomical Society 479, 315 (2018).